# On the inclusion of non-concurrent controls in platform trials with a futility interim analysis

### v1

@@ 5.3 Results
We first examine the marginal and conditional bias, summarized in Figures 3 and 4. Figure 3 shows the marginal bias in the effect estimator for Arm 2 for varying futility bounds $\alpha_{1}$, and under different time trend patterns. The bias is given for the unadjusted model-based estimator, as well as the proposed mean adjusted estimator with $\theta_{1}$ estimated from either both, only period 1, only period 2, or using the Rao-Blackwellized CUMVUE. As discussed in Section 3, the marginal bias of the unadjusted estimator $\tilde{\theta}^{A}_{2}$ reaches its maximum for $\alpha_{1}=0.5$, i.e., in cases where Arm 1 continues in the trial with 50% probability. The mean adjusted estimators provide a substantial reduction in the marginal bias. In particular, the MAEs with $\theta_{1}$ estimated only from period 2 or using the CUMVUE even lead to a conservative effect estimator for Arm 2 and hence to a slightly negative bias. The MAEs with $\hat{\theta}_{1}$ estimated from both periods or only period 1 still maintain some bias due to the inclusion of the period 1 data, which was used for the interim decision. This is more pronounced when the period 1 data are used for the estimation of $\theta_{1}$ exclusively. The obtained bias is consistent across scenarios with no time trend, and linear or stepwise time trends with $\lambda=0.15$. Hence, the mean adjusted estimator is robust to time trends, even if their shape is misspecified relative to the assumption of a stepwise pattern used for the derivation. This holds for all considered operating characteristics. Therefore, in the main paper, we only present results under no time trends, and results for linear time trends. The corresponding results for other time trends are presented in the supplementary material.
@@
Figure 4 shows the bias conditional on Arm 1 continuing after the interim analysis, as a function of the futility bound $\alpha_{1}$, the ratio between sample sizes in periods 1 and 2 $r$, the ratio between sample sizes in Arm 1 and control $a$, and the strength of the time trend $\lambda$. Again, the conditional bias of the model-based estimator is reduced in all cases when using the mean adjusted estimator. This improvement is more pronounced in designs that use a lower futility bound at the interim analysis, since the conditional bias of the unadjusted estimator has the highest value in these cases (see Figure 4-A). The MAEs with $\theta_{1}$ estimated from either both periods or only period 1 still show a positive conditional bias, which also decreases with increasing $\alpha_{1}$. This is because with lower futility bounds, only more extreme samples in period 1, i.e., those with higher estimated treatment effects in Arm 1, will allow for continuation after the interim analysis compared to scenarios with higher futility bounds. This overestimation of $\theta_{1}$ in period 1 leads to lower values of the estimated bias used for the computation of the MAEs, if the period 1 data is included. As a result, there is insufficient bias reduction and hence a remaining positive bias in the treatment effect estimation for Arm 2. On the other hand, the MAEs with $\theta_{1}$ estimated only from period 2 and using the CUMVUE are slightly negatively biased. However, their performance is more robust to the varying $\alpha_{1}$. For increasing sample sizes in period 1, while keeping the period 2 sample sizes constant, the MAE with $\theta_{1}$ estimated only from period 2 becomes more negatively biased, and can even lead to a similar absolute conditional bias as the unadjusted model-based estimator (see Figure 4-B). In this case, the MAE with $\theta_{1}$ estimated using the CUMVUE provides a substantial improvement, since $\hat{\theta}_{1}^{CUMVUE}$ has a lower variance than $\hat{\theta}_{1}$ computed from period 2 only. Moreover, the conditional bias when using the MAEs with $\theta_{1}$ estimated from period 2 or employing the CUMVUE, does not sensitively depend on the allocation ratio between Arm 1 and the control arm (see Figure 4-C), and does not depend on the strength of the time trend $\lambda$ (see Figure 4-D).
@@
Figure 5 shows the conditional root mean squared error (rMSE) of the newly proposed estimator compared to the model-based unadjusted estimator. In all scenarios, the adjusted estimators achieve a lower rMSE than the separate analysis using concurrent controls only. However, in most scenarios, the bias adjusted estimators based on the period 2 estimate of $\theta_{1}$ and the CUMVUE have a larger mean squared error than the unadjusted estimator. For all considered methods, the rMSE is decreasing with increasing sample sizes in period 1 (see Figure 5-B), which is due to the increase in the total sample size in the platform trial in these scenarios.
@@
Next, we assess the properties of the corresponding hypothesis test. The conditional type I error rate is shown in Figure 6. The inflation of the conditional type I error rate caused by applying the unadjusted model-based estimator is substantially reduced when using the mean adjusted estimators. However, the MAEs computed using $\hat{\theta}_{1}$ from period 1 or both periods still show some inflation. The MAEs with $\theta_{1}$ estimated using the CUMVUE yield type I error rates within the simulation error in all considered scenarios. In contrast, the MAE with $\hat{\theta}_{1}$ computed only from period 2 data leads to a strictly conservative test if the sample size in period 1 is much larger than in period 2 (see Figure 6-B). Figure 6-D demonstrates that the type I error rate for given sample sizes and futility bound is not dependent on the strength and pattern of the time trend.
@@
Figure 7 shows the conditional power of the model-based procedures and, for comparison, the separate analysis using only concurrent controls. The highest power is achieved by the unadjusted model-based estimator in all considered cases, which, however, is due to the bias in the treatment effect estimation and inflation in the type I error rate. Nevertheless, the mean adjusted estimators that control the type I error rate (based on $\theta_{1}$ estimates from period 2 and using the CUMVUE) also lead to power gains compared to the separate analysis, as they make use of the non-concurrent control data. As expected, the conditional power increases for increasing period 1 sample sizes (keeping the period 2 sample size fixed) and increasing sample sizes in Arm 1 (see Figures 7-B and 7-C, respectively). As the other operating characteristics, the conditional power does not depend on the strength of the time trend (see Fig. 7-D).
@@
Figure 3: Marginal bias in the treatment effect estimator for Arm 2 for varying futility bound ($\alpha_{1}$) using the unadjusted estimator and the mean adjusted estimator with $\theta_{1}$ estimated from both periods, only period 1, only period 2, or using the CUMVUE. This figure corresponds to a platform trial with sample sizes per arm and period set to 150. The strength of the time trend is $\lambda=0.15$ for both linear and stepwise patterns.
@@
A) Varying futility bound ($\alpha_{1}$). This figure corresponds to a platform trial without time trends, with sample sizes per arm and period set to 150.
B) Varying ratio between the sample sizes in period 1 and period 2 ($r$), where the period 2 sample sizes per arm are fixed. Sample sizes per arm in period 1 are determined by the chosen ratio $r$, i.e., increase or decrease with respect to the period 2 sample sizes $r$-times. This figure corresponds to a platform trial without time trends, with sample sizes per arm in period 2 set to 150, using a futility bound $\alpha_{1}=0.5$.
C) Varying ratio between the sample sizes in Arm 1 and control ($a$), where the control arm sample sizes per period are fixed. Sample sizes in Arm 1 per period are determined by the chosen ratio $a$, i.e., increase or decrease with respect to the control arm sample sizes $a$-times. This figure corresponds to a platform trial without time trends, with sample sizes in the control arm per period set to 150, using a futility bound $\alpha_{1}=0.5$.
D) Varying strength of the time trend ($\lambda$). This figure corresponds to a platform trial with a linear time trend of strength $\lambda$, with sample sizes per arm and period set to 150, using a futility bound $\alpha_{1}=0.5$.
@@
Figure 6: Conditional type I error rate for Arm 2 using the unadjusted estimator $\tilde{\theta}_{2}$ and the mean adjusted estimator $\tilde{\theta}^{A}_{2}$ with $\theta_{1}$ estimated from both periods, only period 1, only period 2, or using the CUMVUE. All figures include a dashed reference line for the nominal significance level of 0.025 and the simulation error is shown as a gray area representing the 95% confidence interval of the simulated type I error rate. Note that in Figure 6-A, this area is not constant over different values of $\alpha_{1}$. This is because for lower values of $\alpha_{1}$, the conditional type I error rate is estimated based on fewer samples, since Arm 1 has a lower probability to continue after the interim analysis in these cases. See the legend of Figure 4 for the explanation of the subfigures A)-D).
@@ 6 Discussion
Existing methods for NCC inclusion have not accounted for interim analyses. We demonstrated that model-based approaches used to adjust for time trends in platform trials can introduce a positive bias and inflate type I error rates once futility interim analyses are included, even in the absence of actual time trends. In particular, conducting an interim analysis in one arm affects the treatment effect estimation in other arms when non-concurrent controls are used. This occurs because the weight of the non-concurrent controls in the model-based estimator depends on the interim result of other arms.
@@
To address this challenge in a platform trial with two experimental arms and two periods, we proposed mean adjusted treatment effect estimators for Arm 2, derived from the original model-based estimator by subtracting its estimated bias. This estimator adjusts for time trends while accounting for the interim analysis in Arm 1. Because the bias-adjusted estimator for Arm 2 depends on an estimate of the treatment effect $\theta_{1}$ of Arm 1, we considered several estimators. In a simulation study, we showed that using the CUMVUE to estimate $\theta_{1}$ minimized the bias of the bias-adjusted estimator of Arm 2 across all considered options. In none of the settings did we observe a positive bias that overestimated the treatment effect.
@@
Moreover, we proposed an adjusted hypothesis test for Arm 2, using the Wald test based on the mean adjusted estimator. This test controls the type 1 error rate when using the CUMVUE to estimate $\theta_{1}$ and yields power gains compared to a separate analysis using only concurrent controls, due to the increased sample size from incorporating non-concurrent controls. However, these power gains are notably smaller than those achieved by the regression model in settings without interim analyses. This is because the mean adjusted estimator accounts simultaneously for time trends and the interim analysis. When using non-concurrent controls, the potential gains in power must be weighed against the added complexity and reduced robustness to model assumptions. The proposed estimator is robust to time trends, even when their shape deviates from the assumed stepwise pattern. However, since the time trend adjustment relies on the assumptions of equal time trends across all arms and additivity of the time trends on the model scale, deviations from these assumptions may lead to bias and inflation of the type I error rate [9].
In this work, we focused on futility interim analyses, where the unadjusted model-based estimator that does not account for the interim analysis is positively biased. In contrast, with interim analyses allowing early stopping for efficacy, the unadjusted model-based estimator is, by symmetry, negatively biased. This occurs because the mean response in the non-concurrent controls is overestimated when Arm 1 continues after the interim analysis, leading to underestimation of the treatment effect in Arm 2. When both futility and efficacy stopping rules are applied, the marginal bias of the model-based estimator is influenced by both mechanisms, and its direction depends on the treatment effect in Arm 1. For small treatment effects in Arm 1, early stopping occurs predominantly for futility, resulting in a positive marginal bias in $\tilde{\theta}_{2}$. For large treatment effects in Arm 1, early stopping occurs predominantly for efficacy, resulting in a negative marginal bias in $\tilde{\theta}_{2}$. For intermediate effect sizes, the opposing impacts of futility and efficacy stopping may partially offset one another, reducing the magnitude of the bias.
@@
If, in addition to the interim analysis for Arm 1, an interim analysis is also performed for Arm 2, this introduces an additional bias in the treatment effect estimate for Arm 2. An efficacy interim analysis for Arm 2 will also impact the type 1 error rate of hypothesis tests, requiring the use of adjusted boundaries.
@@
Extending the proposed methodology to incorporate efficacy stopping, more general timings of interim analyses, trials with more than two experimental arms, and interim analyses for all arms is a topic for future research.
@@
Conflict of Interest Statement The authors declare no potential conflict of interests.

### v2

@@ 5.3 Results
We first examine the marginal and conditional bias, summarized in Figures 4 and 5. Figure 4 shows the marginal bias in the effect estimator for Arm 2 for varying treatment effects in Arm 1, $\theta_{1}$, and under different time trend patterns. The bias is given for the unadjusted model-based estimator, as well as the proposed mean adjusted estimator with $\theta_{1}$ estimated from either both, only period 1, only period 2, or using the Rao-Blackwellized CUMVUE. As discussed in Section 3, the marginal bias of the unadjusted estimator $\tilde{\theta}_{2}$ reaches its maximum in the absolute value for $\theta_{1}=0$ and $\theta_{1}=0.3,$ i.e., in cases where Arm 1 continues in the trial with approx. 50% probability. The mean adjusted estimators provide a substantial reduction in the marginal bias. In particular, this reduction is most pronounced for the MAEs with $\theta_{1}$ estimated only from period 2 or using the CUMVUE, even though they still maintain a slight bias in the extreme cases where Arm 1 continues with approx. 50% probability. The MAEs with $\hat{\theta}_{1}$ estimated from both periods or only period 1 still maintain considerable bias due to the inclusion of the period 1 data, which was used for the interim decision. This is more pronounced when the period 1 data are used for the estimation of $\theta_{1}$ exclusively. The obtained bias is consistent across scenarios with no time trend, and linear, stepwise, seasonal, or random walk time trends with $\lambda=0.15$. Hence, the mean adjusted estimator is robust to time trends regardless of their pattern and strength. This holds for all considered operating characteristics. Therefore, in the main paper, we only present results under no time trends, and results for linear time trends. The corresponding results for other time trends are presented in the supplementary material (Section A).
@@
Figure 5 shows the bias conditional on Arm 1 continuing after the interim analysis, as a function of the treatment effect in Arm 1 $\theta_{1}$, the ratio between sample sizes in periods 1 and 2 $r$, the ratio between sample sizes in Arm 1 and control $a$, and the strength of the time trend $\lambda$. Again, the conditional bias of the model-based estimator is reduced in all cases when using the mean adjusted estimator. This improvement is more pronounced in designs that have more extreme treatment effects in Arm 1, since the conditional bias of the unadjusted estimator has the highest value in these cases (see Figure 5-A). The MAEs with $\theta_{1}$ estimated from either both periods or only period 1 still show some conditional bias, which also decreases as $\theta_{1}$ becomes more moderate. This overestimation or underestimation of $\theta_{1}$ in period 1 leads to lower or higher values of the estimated bias used for the computation of the MAEs, if the period 1 data is included. As a result, there is insufficient bias reduction and hence a remaining positive or negative bias in the treatment effect estimation for Arm 2. The MAEs with $\theta_{1}$ estimated only from period 2 and using the CUMVUE are only very slightly negatively or positively biased. However, their performance is more robust to the varying $\theta_{1}$. Moreover, the conditional bias when using the MAEs with $\theta_{1}$ estimated from period 2 or employing the CUMVUE, does not depend sensitively on the sample size ratio between periods 1 and 2 (see Figure 5-B), or the allocation ratio between Arm 1 and the control arm (see Figure 5-C), and does not depend on the strength of the time trend $\lambda$ (see Figure 5-D).
@@
Figure 6 shows the conditional root mean squared error (rMSE) of the newly proposed estimator compared to the model-based unadjusted estimator. In almost all scenarios, the adjusted estimators achieve a lower rMSE than the separate analysis using concurrent controls only. However, in most scenarios, the bias adjusted estimators based on the period 2 estimate of $\theta_{1}$ and the CUMVUE have a larger mean squared error than the unadjusted estimator. Exceptions are extreme cases, where Arm 1 has a very low probability to continue (see Figure 6-A), where the rMSE of the unadjusted estimator exceeds the rMSE of the mean adjusted estimators due to its very large bias in such settings. For all considered methods that take into account the data from period 1, the rMSE is decreasing with increasing sample sizes in period 1 (see Figure 6-B).
@@
Next, we assess the properties of the corresponding hypothesis test. The conditional type I error rate is shown in Figure 7. The inflation or deflation of the conditional type I error rate caused by applying the unadjusted model-based estimator is substantially reduced when using the mean adjusted estimators. However, the MAEs computed using $\hat{\theta}_{1}$ from period 1 or both periods still show some inflation or deflation. The MAEs with $\theta_{1}$ estimated using the CUMVUE yield type I error rates within the simulation error in all considered scenarios. Figure 7-D demonstrates that the type I error rate for given sample sizes and stopping bounds is not dependent on the strength of the time trend.
@@
Figure 8 shows the conditional power of the model-based procedures and, for comparison, the separate analysis using only concurrent controls. The highest power is achieved by the unadjusted model-based estimator in most of the considered cases, which, however, is due to the bias in the treatment effect estimation and inflation in the type I error rate. Nevertheless, the mean adjusted estimators that control the type I error rate (based on $\theta_{1}$ estimates from period 2 and using the CUMVUE) also lead to power gains compared to the separate analysis, as they make use of the non-concurrent control data. As expected, the conditional power increases for increasing period 1 sample sizes (keeping the period 2 sample size fixed) and increasing sample sizes in Arm 1 (see Figures 8-B and 8-C, respectively). As the other operating characteristics, the conditional power does not depend on the strength of the time trend (see Fig. 8-D).
In the settings considered in this Section, early stopping occurs predominantly for efficacy, which leads to conservative treatment effect estimates and hypothesis tests. In the supplementary material (Section B), we present additional results for settings where only stopping for futility is applied. Such scenarios yield positive bias and type I error rate inflation in the unadjusted estimator, which is considerably reduced by applying the mean adjusted estimators. In particular, the MAE computed using the CUMVUE estimator for $\theta_{1}$ shows the best performance across the considered scenarios.
@@
Figure 4: Marginal bias in the treatment effect estimator for Arm 2 for varying the treatment effect in Arm 1 ($\theta_{1}$) using the unadjusted estimator and the mean adjusted estimator with $\theta_{1}$ estimated from both periods, only period 1, only period 2, or using the CUMVUE. This figure corresponds to a platform trial with sample sizes per arm and period set to 150. The futility boundary was set to $\alpha_{F}=0.5$, while for the efficacy boundary, $\alpha_{E}=0.00264$ was chosen, corresponding to the O’Brien-Fleming boundary for the given sample sizes and $\alpha_{F}$, assuming a significance level of 0.025 in the final analysis. The strength of the time trend is $\lambda=0.15$ for all considered patterns.
@@
A) Varying treatment effect in Arm 1 ($\theta_{1}$). This figure corresponds to a platform trial without time trends, with sample sizes per arm and period set to 150. The futility boundary was set to $\alpha_{F}=0.5$, while for the efficacy boundary, $\alpha_{E}=0.00264$ was chosen, corresponding to the O’Brien-Fleming boundary for the given sample sizes and $\alpha_{F}$, assuming a significance level of 0.025 in the final analysis. Note that we have restricted the range of the values for $\theta_{1}$, such that Arm 1 has a reasonable probability to continue after the interim analysis ($4\%$).
B) Varying ratio between the sample sizes in period 1 and period 2 ($r$), where the period 2 sample sizes per arm are fixed. Sample sizes per arm in period 1 are determined by the chosen ratio $r$, i.e., increase or decrease with respect to the period 2 sample sizes by a factor of $r$. This figure corresponds to a platform trial without time trends, with sample sizes per arm in period 2 set to 150. The futility boundary was set to $\alpha_{F}=0.5$, while the efficacy boundary $\alpha_{E}$ was set in each case to the corresponding O’Brien-Fleming boundary for the given sample sizes and $\alpha_{F}$, assuming a significance level of 0.025 in the final analysis. The treatment effect in Arm 1 was in each case chosen such that a fixed design analysis of Arm 1 would yield approx. 80% power.
C) Varying ratio between the sample sizes in Arm 1 and control ($a$), where the control arm sample sizes per period are fixed. Sample sizes in Arm 1 per period are determined by the chosen ratio $a$, i.e., increase or decrease with respect to the control arm sample sizes by a factor of $a$. This figure corresponds to a platform trial without time trends, with sample sizes in the control arm per period set to 150. The futility boundary was set to $\alpha_{F}=0.5$, while for the efficacy boundary, $\alpha_{E}=0.00264$ was chosen, corresponding to the O’Brien-Fleming boundary for the given sample sizes and $\alpha_{F}$, assuming a significance level of 0.025 in the final analysis. The treatment effect in Arm 1 was in each case chosen such that a fixed design analysis of Arm 1 would yield approx. 80% power.
D) Varying strength of the time trend ($\lambda$). This figure corresponds to a platform trial with a linear time trend of strength $\lambda$, with sample sizes per arm and period set to 150. The futility boundary was set to $\alpha_{F}=0.5$, while for the efficacy boundary, $\alpha_{E}=0.00264$ was chosen, corresponding to the O’Brien-Fleming boundary for the given sample sizes and $\alpha_{F}$, assuming a significance level of 0.025 in the final analysis. The treatment effect in Arm 1 was in each case chosen such that a fixed design analysis of Arm 1 would yield approx. 80% power.
@@
Figure 7: Conditional type I error rate for Arm 2 using the unadjusted estimator $\tilde{\theta}_{2}$ and the mean adjusted estimator $\tilde{\theta}^{A}_{2}$ with $\theta_{1}$ estimated from both periods, only period 1, only period 2, or using the CUMVUE. All figures include a dashed reference line for the nominal significance level of 0.025 and the simulation error is shown as a gray area representing the 95% confidence interval of the simulated type I error rate. Note that in Figure 7-A, this area is not constant over different values of $\theta_{1}$. This is because for more extreme values of $\theta_{1}$, the conditional type I error rate is estimated based on fewer samples, since Arm 1 has a lower probability to continue after the interim analysis in these cases. See the legend of Figure 5 for the explanation of the subfigures A)-D).
@@ 6 Discussion
Existing methods for NCC inclusion have not accounted for interim analyses. We demonstrated that model-based approaches used to adjust for time trends in platform trials can introduce a positive bias and inflate type I error rates once futility interim analyses are included, even in the absence of actual time trends. Similarly, efficacy stopping may lead to negative bias and deflation of the type I error rate. In this work, we showed that conducting an interim analysis in one arm affects the treatment effect estimation in other arms when non-concurrent controls are used. This occurs because the weight of the non-concurrent controls in the model-based estimator depends on the interim result of other arms.
From the analytical expression for the bias, we can see that inflated or deflated marginal type I error rates of the unadjusted model-based estimator for Arm 2 are most pronounced when (i) borrowing from NCC is substantial (i.e., large $\varrho$), (ii) the interim decision for Arm 1 has a moderate probability of stopping or continuing (i.e., continuation probability of around 50%), and (iii) period 1 and period 2 sample sizes are similar. Conversely, the inflation/deflation is negligible when NCC borrowing is minimal (i.e., $\varrho\approx 0$, for example when there is a very large sample size for the concurrent control in period 2), when the interim rule is extreme (meaning that Arm 1 almost always stops or continues), or when periods 1 and 2 substantially differ in sample sizes.
@@
To address this challenge in a platform trial with two experimental arms and two periods, we proposed mean adjusted treatment effect estimators for Arm 2, derived from the original model-based estimator by subtracting its estimated bias. This estimator adjusts for time trends while accounting for the interim analysis in Arm 1. Because the bias-adjusted estimator for Arm 2 depends on an estimate of the treatment effect $\theta_{1}$ of Arm 1, we considered several estimators. In a simulation study, we showed that using the CUMVUE to estimate $\theta_{1}$ minimized the bias of the bias-adjusted estimator of Arm 2 across all considered options.
@@
Moreover, we proposed an adjusted hypothesis test for Arm 2, using the Wald test based on the mean adjusted estimator. This test controls the type I error rate in the considered scenarios when using the CUMVUE to estimate $\theta_{1}$ and yields power gains compared to a separate analysis using only concurrent controls, due to the increased sample size from incorporating non-concurrent controls. However, these power gains are notably smaller than those achieved by the regression model in settings without interim analyses. This is because the mean adjusted estimator accounts simultaneously for time trends and the interim analysis. When using non-concurrent controls, the potential gains in power must be weighed against the added complexity and reduced robustness to model assumptions. The proposed estimator is robust to time trends regardless of their pattern. However, since the time trend adjustment relies on the assumptions of equal time trends across all arms and additivity of the time trends on the model scale, deviations from these assumptions may lead to bias and inflation of the type I error rate [3].
@@
If, in addition to the interim analysis for Arm 1, an interim analysis is also performed for Arm 2, this introduces an additional bias in the treatment effect estimate for Arm 2. An efficacy interim analysis for Arm 2 will also impact the type I error rate of hypothesis tests, requiring the use of adjusted boundaries.
@@
Recent methodological work has identified a related issue in platform trials that arises when allowing for a change in the control arm. For instance, when an experimental treatment is found to be superior and becomes the new standard of care. In this setting, a decision must be made whether to keep or discard the information collected before the control is changed. Greenstreet et al. show that this choice can also substantially alter the conditional type I error rate and conditional power [14].
In this work, we assumed that the final analysis of both experimental arms is performed at the same time. The resulting 2-period design represents an extreme case with respect to the bias of the treatment effect estimator for Arm 2, because it maximizes the overlap between the two experimental arms. A larger overlap increases the contribution of non-concurrent controls in the estimator from the period-adjusted regression model, whereas no overlap implies no borrowing. As a result, the interim decision on Arm 1 has a stronger impact on the Arm 2 estimator in this setting. If Arm 1 has an earlier final analysis, the trial includes an additional third period with only Arm 2 and control, thereby reducing the overlap between the two experimental arms.
Extending the proposed methodology to incorporate more general timings of interim and final analyses, trials with more than two experimental arms, and interim analyses for all arms is a topic for future research. These extensions would require accounting for additional complexity in how NCC data contributes to the estimator for late-entering arms. In particular, if interim or final analyses are conducted when no additional arms are entering the trial, the trial is split into more periods. Each additional period introduces a new period effect in the regression model, and the weight assigned to the NCC data is determined by the amount of overlapping sample size between experimental arms. Moreover, in platform trials with more than two experimental arms, the treatment effect estimator for late-entering arms may depend on interim decisions of multiple earlier arms. This means that the bias correction becomes conditional on multiple continuation events rather than a single interim outcome. In such complex cases, closed-form bias adjustments may no longer be tractable, and simulation-based estimation for the bias and variance may be required.
@@
Martin Posch is a member of RealiseD, which is supported by the Innovative Health Initiative Joint Undertaking (IHIJU) under grant agreement No 101165912. The JU receives support from the European Union’s Horizon Europe research and innovation programme and COCIR, EFPIA, EuropaBío, MedTech Europe, and Vaccines Europe. Views and opinions expressed are those of the authors only. They do not necessarily reflect those of the Innovative Health Initiative Joint Undertaking and its members, EMA, the Norwegian Medical Products Agency or any pharmaceutical company.
Conflict of Interest Statement The authors declare no potential conflict of interest.